### v2

@@ Appendix B Hyper-parameter optimization
Appendix C Loss curves
Figure 8 reports the loss curves as a function of epochs and wall-clock time for the different algorithms, corresponding to the accuracies reported in the main text.
Figure 8: Performance of BP and DBP using $\kappa=0$ (decorrelation) and $\kappa=0.5$ (whitening). Reported results are the average of three randomly initialized networks, where minimal and maximal value are indicated by the error bars. a) Train loss as a function of the number of epochs. b) Test loss as a function of the number of epochs. c) Train loss as a function of wall-clock time. d) Test loss as a function of wall-clock time.
Appendix D Train versus test loss comparison
Figure 9 shows the train versus test losses for a ResNet-18 model trained on ImageNet for both BP and DBP. As DBP’s test loss is slightly lower than the BP test loss at the same value of the train loss, we conclude that DBP yields slightly better generalization performance compared to BP.
Figure 9: Scatterplot of train versus test loss for BP and DBP.
Appendix E Peak performance comparison
Table 1 provides an overview of peak performance for the different algorithms, as well as the epoch number and time at which peak test performance was achieved.
Table 1: Peak performance for BP, DBP in terms of loss and accuracy. Epoch and Time indicate the epoch number and wall-clock time at which peak accuracy was achieved. Best results shown in boldface, where we focus on test results as this quantifies generalization performance.
Loss Accuracy Epoch Time (hours)
Method $\kappa$ Train Test Train Test Train Test Train Test
BP - 1.24 2.90 99.4 54.1 49 12 16.0 3.7
DBP 0 1.21 2.86 99.5 54.8 50 7 18.4 2.3
DBP 0.5 1.22 2.84 99.5 55.2 50 7 18.4 2.3
Appendix F Impact of $\kappa$ parameter on CIFAR10 performance
Figure 10 shows train and test performance for a small three-layer ConvNet trained on CIFAR10 for BP and for several settings of DBP where the $\kappa$ parameter was varied. Setting $\kappa=0$ reached the highest peak test accuracy for this particular task. Performance gains for this smaller model are even more pronounced yet result in a smaller reduction in carbon consumption due to the model’s smaller size. Note further that peak test performance for DBP is larger than that of BP, again indicating improved generalization performance. A batch size of 256 and learning rate of $1\times 10^{-3}$ were used in all experiments. The decorrelation/whitening learning rate was also set to $1\times 10^{-3}$.

### v4

@@ Appendix C Hyper-parameter optimization
Appendix D Impact of whitening for ConvNets on CIFAR-10
To investigate the effect of the $\kappa$ parameter in a different setting than large (residual) DNNs, a small three-layer ConvNet was trained on the CIFAR-10 [Krizhevsky, 2009] dataset. The model was trained using BP as welll as DBP under different values of the $\kappa$ parameter. As shown in Figure 10, a setting of $\kappa=0$ reached the highest peak test accuracy, though differences were small for for values less than $\kappa=0.6$. Note further that performance gains for this smaller model are even more pronounced than for the larger models.